OpenStudy (anonymous):

Select the equation of the line that passes through the point (2, 6) and is perpendicular to the line x = 4.

OpenStudy (anonymous):

The line x = 4 has an undefined slope. Basically, this line is absolutely vertical on a Cartesian grid. So if we're looking for a another line perpendicular to x = 4, we'll have to use a horizontal line (a line with a slope of m = 0 on a Cartesian grid.) First, let's draw out the point-slope format: \[y = mx + b\] Next, we plug in our point (2, 6). \[6 = m \left( 2 \right) + b\] \[6 = 0\left( 2 \right) + b\] \[6 = b\] And since we're looking for a horizontal line, we're left with: \[y = 6\]

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

can you help me with one more please @Cardinal_Carlo

OpenStudy (anonymous):

Sure, just ask away.

OpenStudy (anonymous):

Oh, but close this question first.

OpenStudy (anonymous):

Choose the equation of the vertical line passing through the point (−2, 3).

OpenStudy (anonymous):

@Cardinal_Carlo